# Global Non-Axisymmetric Hall Instabilities in a Rotating Plasma

### v1

@@ 2.4 Over-reflection of $k=0$ modes in azimuthal field
Figure 5 plots the function $Q(\omega,s)$ corresponding to the fastest growing mode attained by this system. Also plotted is the structure of eigenmode. Note that the ‘potential well’ in this case is relatively shallow. The phase of the eigenmode fortuitously illustrates the transition between the imaginary being dominant in the region left of the first turning point, and the real part dominant to the right of the second turning point. This is what one qualitatively expects by applying asymptotic (WKB) turning point theory. As shown in appendix A, the solutions may be approximated by parabolic cylinder functions that are oscillatory in the region outside of the shaded area, and behave as growing and decaying exponential functions inside the shaded region. Extending the discussion of wave over-reflection generating global modes in non-conducting fluid disk model to the present EMHD whistler model in a manner similar to R. Narayan et al. (1987) is conceptually straightforward. The technical details and analysis of equation (12) are discussed in appendix A.
3 Numerical results for Hall-MHD
In this section we provide numerical calculations of the eigenvalues of the incompressible Hall-MHD system in a cylinder using the same technique described in section 2.2 for EMHD. Although we do not include resistivity, we note that modes with large $|\boldsymbol{k}|$ are stabilized by dissipation (F. Ebrahimi et al., 2011; F. Ebrahimi & A. Haywood, 2025; A. Haywood & F. Ebrahimi, 2025b). In this section we focus on the most global (i.e. lowest $m,k$) modes, since these modes would be the least affected by dissipative mechanisms. To assess the effect of varying strength of the Hall term, we compare the results for different values of $d_{i}$ to both the MHD result, and the EMHD limit with $d_{i}=1$ for both azimuthal and vertical magnetic field configurations.
We compare a subset of the eigenvalue results with linear initial value calculations using the extended MHD code NIMROD (C. Sovinec et al., 2004). NIMROD (non-ideal magnetohydrodynamics with rotation, open discussion), uses high-order spectral element method to solve both linear and nonlinear extended MHD. The domain is discretized using a 2-D grid of high-order polynomial finite elements in the rzr-z plane, and a pseudospectral collocation method with a Fourier basis for the periodic azimuthal coordinate, $\phi$. The equations are integrated in time using a semi-implicit leapfrog time-stepping scheme. The accuracy and numerical stability of the two-fluid model in NIMROD has been successfully benchmarked with simulations of linear and nonlinear non-ideal MHD instabilities (C. R. Sovinec & J. R. King, 2010).
We consider the situation where the plasma is enclosed between two conducting cylindrical walls. In this case, the appropriate boundary conditions applicable to equations (4)–(7) are $\psi=0$ and $w=0$ at the inner and outer wall locations. In NIMROD, and in the eigenvalue code, the equations are solved for $\boldsymbol{v}$ and $\boldsymbol{b}$, and the normal components are set to zero at the inner and outer wall.
We first consider the case in which the magnetic field is oriented vertically. The linear growth of the $m=0$ modes in Hall-MHD has been studied (M. Wardle, 1999; F. Ebrahimi et al., 2011; M. Wardle & R. Salmeron, 2012). Considering non-axisymmetric modes, we find that it is generally the case that the $m=0$ is the fastest growing mode at low magnetic field strengths. For completeness, we compare the growth rates of the $m=0$ to the $m=1$ modes at two values of $d_{i}$ as a function of $V_{A,z}/r_{1}\Omega_{0}$ in figure 6.
For both values of $d_{i}$ the $m=0$ has a larger growth rate when $|B_{z}|$ is small, and a higher maximum growth rate. There is an intermediate value of $|B_{z}|$, however, where the $m=0$ mode is stable, and the $m=1$ is the dominant mode. We do not plot the mode frequencies here, since the $m=0$ has $\omega_{r}=0$, and the associated frequency for the $m=1$ modes are shown in figure 7.
Fig. 7 plots the growth rates and frequencies of the fastest growing modes as a function of the magnetic field strength for the MHD model, three values of $d_{i}$ in the Hall-MHD model, and the result from EMHD. The effect of the Hall term is significant even at modest values of $d_{i}$. At $d_{i}=0.1r_{1}$, the ion skin depth is only 2.5% of the width of the cylinder, yet the growth rates of the fastest growing modes are more than twice as fast as in the MHD case. The Hall term destabilizes a different high-frequency ($\omega\sim\Omega$) branch of non-axisymmetric instability that grows significantly faster than either the $m=1$ MRI or the low frequency global curvature mode in the MHD limit (F. Ebrahimi & M. Pharr, 2022). These are the whistler modes that we found in the EMHD limit in section 2.
Figure 7: Growth rates and frequencies of the most unstable mode with $m=1,k=\pi/4$ as a function of the vertical magnetic field strength measured by the inverse Alfvénic Mach number. We show the MHD result, Hall MHD with two values of $d_{i}$, and the EMHD limit with $d_{i}=1.0$, for comparison. As the strength of the Hall effect ($d_{i}$) increases, the growth rate develops an asymmetry with respect to the direction magnetic field.
As $d_{i}$ is increased, the dispersion relation develops an asymmetric dependence on direction of $\boldsymbol{B}$ relative to the axis of rotation as the system transitions from the MHD to the Hall-dominant regime. In the right-handed cylindrical coordinates used, we assume $\boldsymbol{\Omega}\propto\hat{\mathbf{z}}$, so that positive $B_{z}$ corresponds to a magnetic field parallel to the axis of rotation. As $d_{i}$ increases, the $m=1,k=1$ modes with $B_{z}<0$ have lower growth rates than the modes with $B_{z}>0$, but remain unstable at larger values of $|B_{z}|$.
The radial profiles of the eigenfunctions in the Hall-MHD calculations for $B_{z}>0$ further suggest that the mode observed here is the whistler mode seen in the EMHD model. In the full Hall-MHD system, however, the mode structure is modified by coupling between the electron and ion flow. Fig. 8 shows the radial profiles of the solutions to the Hall-MHD eigenvalue problem at $d_{i}=0.1$ for $V_{A,z}=0.1r_{1}\Omega_{1}$ and $V_{A,z}=0.3r_{1}\Omega_{0}$. Similar to the EMHD case, as the strength of the vertical field decreases, the solutions become increasingly oscillatory. This shortening of the radial wavelength is expected from the whistler modes, since, based on equation (12), the effective radial wave number, $\delta$ is proportional to $\bar{\omega}/kd_{i}V_{A,z}$. Since $\bar{\omega}$ is $\mathcal{O}(1)$ for unstable modes, as $V_{A,z}\rightarrow 0$, $|\delta|$ increases.
Figure 8: Radial structure of eigenfunctions for Hall-MHD system with $d_{i}=0.1r_{1}$ and vertical magnetic fields $V_{A,z}=0.1r_{1}\Omega_{0}$ (left), $V_{A,z}=0.3r_{1}\Omega_{0}$ (center) and $V_{A,z}=0.5r_{1}\Omega_{0}$ (right)
This argument can also qualitatively account for the dependence of the growth rate on $d_{i}$. As $d_{i}$ increases, smaller values of $V_{A,z}$ correspond to the same value of $|\delta|$ for a given axial wave number, $k$.
When the axial field is anti-parallel with the axis of rotation, $B_{z}<0$, the whistler modes are not present for $m=1,k=\pi/4$. Instead, the most unstable modes are more similar in radial structure to the low-frequency global ‘curvature modes’ of F. Ebrahimi & M. Pharr (2022) that have been modified by the Hall term. As $d_{i}$ increases, the growth rates of the curvature modes for $B_{z}<0$ decrease. Figure 9 shows the eigenmodes from the same three $|V_{A,z}|$ as in figure 8, but with $V_{A,z}<0$.
Figure 9: Radial structure of eigenfunctions for Hall-MHD system with $d_{i}=0.1r_{1}$ and vertical magnetic fields $V_{A,z}=-0.1r_{1}\Omega_{0}$ (left), $V_{A,z}=-0.3r_{1}\Omega_{0}$ (center) and $V_{A,z}=-0.5r_{1}\Omega_{0}$ (right)
When the field is purely azimuthal, $B_{\phi}\neq 0,B_{z}=0$, the asymmetric effect of the Hall term is more pronounced. Fig. 10 plots the growth rates and frequencies of the fastest growing $m=1,k=\pi/4$ mode as a function of the azimuthal field strength. When $B_{\phi}>0$ we see a significant increase in the growth rates of the most unstable non-axisymmetric modes with $k>0$ as $d_{i}$ is increased. With increasing $d_{i}$, the range of $B_{\phi}$ over which the modes are unstable is reduced. This is also consistent with the results of the EMHD model. Figure 10 only shows $k>0$ modes, but we note that when $k<0$ the dispersion relation is mirrored about the vertical axis. When $B_{\phi}<0$ the same growth rates are observed for modes with $k<0$ as for the $k>0$ modes when $B_{\phi}>0$.
Figure 10: Growth rates and frequencies of the most unstable mode with $m=1,k=\pi/4$ as a function of the azimuthal magnetic field strength at the inner boundary (in units of the inverse Alfvénic Mach number). The MHD result, Hall-MHD(HMHD) with two values of $d_{i}$, and the EMHD limit at $d_{i}=1$ are shown. When $k=-\pi/4$, the HMHD and EMHD points are mirrored about the vertical axis.
In contrast to pure EMHD limit, the Hall-MHD system has additional unstable modes with $k>0$ when $B_{\phi}<0$, and $k<0$ when $B_{\phi}>0$ that display identical behavior as $|B_{\phi}|$ increases. These modes do not have a counterpart in EMHD, and they exist at increasingly larger values of $|B_{\phi}|$ at larger $d_{i}$. In particular, we note that the maximum value of the growth rate in the $d_{i}=1.0$ case in fig. 10 occurs at $|V_{A,\phi}|=2r_{1}\Omega_{0}$. These modes are eventually stabilized by sufficiently strong magnetic fields. In the $d_{i}=0.1$ case, this occurs near $|V_{A,\phi}|\approx 1.5r_{1}\Omega_{0}$, and in the $d_{i}=1.0$ case, the unstable parameter range extends nearly to $|V_{A,\phi}|\approx 3r_{1}\Omega_{0}$. Based on the behavior of the mode frequencies, we surmise that there are at least two distinct modes present in the calculation. As $|B_{\phi}|\rightarrow 0$, a higher frequency mode with very slow growth dominates.
Figure 11: Radial profiles of the $m=1,k=\pi/4$ HMHD eigenfunctions with $d_{i}=0.1$ with azimuthal fields $V_{A,\phi}=-0.1\Omega_{0}r_{1}$ (left), and $V_{A,\phi}=-0.5\Omega_{0}r_{1}$ (center), and $V_{A,\phi}=-1\Omega_{0}r_{1}$.
As $|B_{\phi}|$ increases, a the lower frequency mode is destabilized. Unlike the whistler and MRI-type modes, the frequency of this mode is effectively constant as $|B_{\phi}|$ varies.
The polarization of the HMHD ion cyclotron wave in the large $d_{i}$ limit is such that the perturbed flow scales as $|\boldsymbol{v}|\sim kd_{i}V_{A}$, and the magnetic field as $|\boldsymbol{b}|\sim\mathcal{O}(1)$. Figure 13 plots the profiles of $v_{r}$ and $b_{r}$ from the most unstable mode at $V_{A,\phi}=-2\Omega_{0}r_{1}$. The magnitudes of the perturbed $\boldsymbol{v}$ and $\boldsymbol{b}$ approximately equal, which is expected for the polarization of the ion-cyclotron mode at $d_{i}=1$.
Figure 13: Radial profile of the $m=1,k=\pi/4$ eigenfunction of the Hall-MHD system with $d_{i}=1r_{1}$ at $V_{A,\phi}=-2\Omega_{0}r_{1}$
To complement and verify the results of the numerical eigenvalue calculations, we compare the results for the $d_{i}=1$ case to linear initial value calculations using the Hall-MHD model in NIMROD. Small amplitude fluctuations are initialized in for a single azimuthal mode number, and the linear equations are evolved in time until a dominant unstable mode emerges. The linear growth rate is computed by calculating the change in perturbed magnetic and kinetic energies between successive time steps.
Figure 14 plots the growth rates from NIMROD calculations in at a few values of azimuthal field strength for $d_{i}=1r_{1}$. The growth rates agree at $V_{A,\phi}/r_{1}\Omega_{0}=2$, but NIMROD reports significantly higher growth for modes at lower magnetic field strength. This discrepancy owes to the fact that both the axial and radial dimensions in NIMROD are discretized using high-order finite elements. As a result, many axial wave numbers are represented by the numerical grid. The calculation of the growth rate captures the axial wavenumber with the largest growth rate, which; for low magnetic fields, increases with $k$.
Figure 14: Growth rate and frequency of the $m=1,k=\pm\pi/4$ mode from the eigenvalue code, and growth rates from linear NIMROD calculations.
At larger magnetic field strength, however, the NIMROD calculation finds that the fastest growing mode is the same $m=1,k=\pm\pi/4$ that we calculated using the eigenvalue code. Figure 15 plots the radial component of the $m=1$ velocity and magnetic field perturbations from the NIMROD calculation with $V_{A,\phi}/r_{1}\Omega_{0}=1$ and $V_{A,\phi}/r_{1}\Omega_{0}=2$.
Figure 15: Contours of $v_{r}$ and $b_{r}$ from NIMROD calculations at $d_{i}=1.0$ for two different azimuthal field strengths. $V_{A,\phi}=r_{1}\Omega$ (left) and $V_{A,\phi}=2r_{1}\Omega_{0}$ (right). The vertical mode number, $k$, is dramatically different between the two cases.
The faster growing, high-$k$ mode observed when $V_{A,\phi}/r_{1}\Omega_{0}=1$ is radially localized near the point of maximum flow shear in the domain. At the larger field strength, the low-$k$, global mode is the most unstable.
4 Conclusion
@@
We have shown that Hall-MHD leads to additional global non-axisymmetric instabilities in a differentially rotating plasma. The Hall EMF breaks the degeneracy of the linear incompressible MHD dispersion relation, yielding two separate branches that, in the large $|\boldsymbol{k}|d_{i}$ limit, become whistler and ion-cyclotron waves. Both of these waves are destabilized by the flow shear. The coupling of the Hall-MHD waves to the background flow shear leads to new branches of global instabilities. The most interesting features of these global, non-axisymmetric modes are that they grow much faster than non-axisymmetric modes in the MHD model, and that they are unstable at larger field strengths.
By appealing to the EMHD limit, where the dynamical coupling between the electrons and the ions is weak, we identified a subset of modes as whistler waves that are driven unstable by the flow shear. The local analysis and stability criteria derived from the EMHD are a generalization of previous discussions of the Hall-shear or magneto-shear instabilities (M. W. Kunz, 2008; C. Bejarano et al., 2011) to the non-axisymmetric case, including the effect of geometric curvature. A subset of the unstable whistler modes with $k=0$ in the azimuthal field case are shown to be destabilized by a ‘corotation amplifier’ mechanism similar to those discussed by R. S. Lindzen & K. K. Tung (1978), R. Narayan et al. (1987), and D. Tsang & D. Lai (2008). Crucially, modes of this type are not predicted by local dispersion relations.
@@
In the full Hall-MHD system, we varied the strength of the hall term by adjusting $d_{i}$. When the Hall effect is weak, $d_{i}/r_{1}\ll 1$, the behavior of the $m=1,k=\pi/4$ mode is similar to the ideal MHD limit in the vertical field configuration. When $d_{i}/r_{1}=1$, the fastest growing modes are unstable whistler waves when $\boldsymbol{B}\cdot\boldsymbol{\Omega}>0$. The frequencies and growth rates are similar to those of the EMHD model. When the vertical field is opposite the axis of rotation, ($\boldsymbol{B}\cdot\boldsymbol{\Omega}<0$), the unstable modes are slightly a slower growing version of the global mode observed in the MHD case. (F. Ebrahimi & M. Pharr, 2022) The intermediate case, where $d_{i}/r_{1}=0.1$, the effect of the Hall term is more subtle. In particular, the Hall-MHD calculations with purely azimuthal field show large-scale, global instabilities that are identified as ion-cyclotron waves. These modes are not described by the EMHD limit. As the ion skin depth is increased, the system supports non-axisymmetric instabilities at increasingly larger magnetic field strengths. At $d_{i}/r_{1}=1$ the whistler mode branch of the HMHD model agrees with the results of the EMHD model. The range of ion-cyclotron wave instability shifts towards even larger $|B|$. At high $d_{i}/r_{1}$, these results have also been verified by initial value NIMROD code, where the same growth rates and similar radial mode structures where obtained.
@@
The Hall term modifies the linear mode spectrum of non-axisymmetric modes at relatively modest values of the normalized ion skin-depth. We have normalized $d_{i}$ in terms of the radius of the inner wall of the annulus, but it is more common to discuss how it compares to the disk scale height. In our EMHD and Hall-MHD calculations, the height of the disk is $\ell=8r_{1}$. The values of $d_{i}/\ell$ considered are $1.25\times 10^{-3},1.25\times 10^{-2}$ and $1.25\times 10^{-1}$. We compare this to estimates for protoplanetary disks used by M. W. Kunz & G. Lesur (2013), where the relevant parameter is $\ell_{H}\equiv\sqrt{(\rho/\rho_{i})}d_{i}$. Based on a minimum-mass solar nebula model, W. Béthune et al. (2016) gives and estimate of the range of $\ell_{H}/h$, where $h$ is the scale height of the disk (equivalent to $\ell/2$ in our case), to be between $10^{-2}$, and $10^{2}$. Thus, even though Ohmic dissipation and ambipolar diffusion are neglected, the values of $d_{i}$ and the corresponding low-frequency global Hall-MHD modes obtained here remain relevant to protoplanetary disks.
@@
Subsequent work will address how the more virulent non-axisymmetric linear instabilities in the Hall-MHD system may affect nonlinear momentum transport. Global nonlinear Hall-MHD NIMROD calculations are ongoing that aim to quantify the level of momentum transport that can be attributed to the global, non-axisymmetric modes. Concerning linear theory – since we are interested in global stability of the differentially rotating system, we aim to extend these results by considering how they are affected by additional inhomogeneities in the steady-state (radial and vertical stratification, magnetic geometry, etc.). It is likely that a simplified description similar to the EMHD model could also be derived for the ion-cyclotron modes in the large $kd_{i}$ limit.
@@ A.1 One turning point case
In the case there is only one zero of $Q$ in the domain, there is a single turning point. Denote the zero of $Q$ by $\mu$. Then, in this case, the solutions are exponential in the region $1\leq s\leq\mu$, and oscillatory in the region $s_{0}\leq s\leq r_{2}/r_{1}$. Using the technique of R. E. Langer (1959) we can approximate the solution in this case as
@@ A.2 Two turning point case
When $V_{A}\ll 1$, there are two roots of $Q$ located at $\left(\frac{m}{2\omega}\left(1+\sqrt{1\pm 4\omega d_{i}V_{A}\sqrt{m^{2}-1/4}}\right)\right)^{2/3}$. In this case, the domain is divided into two regions where the solution is wavelike, $Q>0$ separated by an ‘evanescent region’ region where the solution is damped, $Q<0$. Let $\mu_{1},\mu_{2}$ denote the locations of the two zeroes of $Q$, and assume $Q<0$ for $\mu_{1}\leq s\leq\mu_{2}$. Then, it can be shown using a Liouville-Green transformation that the solutions are asymptotic to parabolic cylinder functions  (A. H. Nayfeh, 2000). In particular, define $\zeta(s)$ such that when $\mu_{1}\leq s\leq\mu_{2}$
@@
$a$ can be interpreted as the size of the forbidden region in Similar expressions involving suitably chosen branches of the square roots ensure $\zeta(s)$ is a regular function, and that the regions where $Q>0$ are mapped to $\zeta^{2}>1$, and the region where $Q<0$ is mapped to the region $\zeta^{2}<1$. Namely, for $1\leq s\leq\mu_{1}$, set
@@
where $E,E^{*}$ are the complex solutions given by M. Abramowitz & I. A. Stegun (1974). As noted by R. Narayan et al. (1987), the functions $E$ and $E^{*}$ represent radially outward and inward propagating waves, respectively. The boundary conditions on $u$ are transformed into the conditions

### v2

@@ III.4 Over-reflection of $k=0$ modes in azimuthal field
Figure 16 plots the function $Q(\omega,s)$ corresponding to the fastest growing mode attained by this system. Also plotted is the structure of eigenmode. Note that the ‘potential well’ in this case is relatively shallow. The phase of the eigenmode fortuitously illustrates the transition between the imaginary being dominant in the region left of the first turning point, and the real part dominant to the right of the second turning point. This is what one qualitatively expects by applying asymptotic (WKB) turning point theory. As shown in appendix A, the solutions may be approximated by parabolic cylinder functions that are oscillatory in the region outside of the shaded area, and behave as growing and decaying exponential functions inside the shaded region. Extending the discussion of wave over-reflection generating global modes in non-conducting fluid disk model to the present EMHD whistler model in a manner similar to Narayan et al. (1987) is conceptually straightforward. The technical details and analysis of equation (19) are discussed in appendix A.
IV Conclusion
@@
We have investigated the stability of rotating plasmas in regimes beyond MHD using analytical and numerical analysis via both eigenvalue and initial value calculations. We find that Hall-MHD leads to additional global non-axisymmetric instabilities in a differentially rotating plasma. The Hall EMF breaks the degeneracy of the linear incompressible MHD dispersion relation, yielding two separate branches that, in the large $|\boldsymbol{k}|d_{i}$ limit, become whistler and ion-cyclotron waves. Both of these waves can be destabilized by the flow shear. The coupling of the Hall-MHD waves to the background flow shear leads to new branches of global instabilities. In both magnetic geometries – axial and azimuthal – increasing the magnetic field strength preferentially stabilizes small wavelength modes, allowing the large-scale, global modes to dominate. We make two interesting observations about these global, non-axisymmetric Hall-MHD modes. First, there are some parameters at which they grow much faster than non-axisymmetric modes in the MHD model. The second is that these modes may be unstable at much larger field strengths than any instabilities in the MHD model.
@@
In the full Hall-MHD system, we varied the strength of the hall term by adjusting $d_{i}/r_{1}$. When $d_{i}/r_{1}=1$, the fastest growing non-axisymmetric modes are unstable whistler waves at weak magnetic field. This is confirmed through comparison with the results from the EMHD model. At higher field strengths, the whistler modes are stabilized, and global, long-wavelength ‘ion-cyclotron’ modes are the most unstable. As the ion skin depth is increased, the system supports non-axisymmetric instabilities at increasingly larger magnetic field strengths. The range of ion-cyclotron instability in the azimuthal field case, in particular, moves towards values of $|B_{\phi}|$ corresponding to sub-Alfvénic flow. This general observation is consistent with the notion that the Hall effect decouples the ion motion from the magnetic field, so that the stabilizing contribution from field-line bending is weaker than in the MHD model. The numerical results from CYL_SPEC have been verified by comparison with initial value calculations using the NIMROD code, where the same growth rates and similar radial mode structures where obtained.
By appealing to the EMHD limit, where the dynamical coupling between the electrons and the ions is weak, we studied the subset of HMHD modes corresponding to whistler waves that are driven unstable by the flow shear. The local analysis and stability criteria derived from the EMHD are a generalization of previous discussions of the Hall-shear or magneto-shear instabilities (Kunz, 2008; Bejarano et al., 2011) to the non-axisymmetric case in a rotating flow, including the effect of geometric curvature. The numerical calculations interrogated the dependence of the whistler modes on the aspect ratio of the disk geometry, finding that lower aspect ratios support growing modes at larger field strengths. Both local and global stability criteria for the EMHD system are presented.
A subset of the unstable whistler modes with $k=0$ in the azimuthal field case are shown to be destabilized by a ‘corotation amplifier’ mechanism similar to those discussed by Lindzen and Tung (1978), Narayan et al. (1987), and Tsang and Lai (2008). Crucially, unstable modes of this type are not predicted by local dispersion relations, and have not been previously identified.
@@
The Hall term modifies the linear mode spectrum of non-axisymmetric modes at relatively modest values of the normalized ion skin-depth. We have normalized $d_{i}$ in terms of the radius of the inner wall of the annulus, but it is more common to discuss how it compares to the disk scale height. In our EMHD and Hall-MHD calculations, the height of the domain is $\ell=8r_{1}$. The values of $d_{i}/\ell$ considered are $1.25\times 10^{-3},1.25\times 10^{-2}$ and $1.25\times 10^{-1}$. We compare this to estimates for protoplanetary disks used by Kunz and Lesur (2013), where the relevant parameter is $\ell_{H}\equiv\sqrt{(\rho/\rho_{i})}d_{i}$. Based on a minimum-mass solar nebula, Béthune et al. (2016) gives an estimate of the range of $\ell_{H}/h$, where $h$ is the scale height of the disk, between $10^{-2}$ and $10^{2}$. Thus, the values of $d_{i}$ and the corresponding low-frequency global Hall-MHD modes obtained here may be relevant to protoplanetary disks.
@@
Subsequent work will address how the more virulent non-axisymmetric linear instabilities in the Hall-MHD system may affect nonlinear momentum transport. Global nonlinear Hall-MHD NIMROD calculations are ongoing that aim to quantify the level of momentum transport that can be attributed to the global, non-axisymmetric modes. Concerning linear theory – since we are interested in global stability of the differentially rotating system, we aim to extend these results by considering how they are affected by inhomogeneities in the steady-state (radial and axial stratification, gradients in the magnetic field, etc.).
@@ A.1 One turning point case
In the case there is only one zero of $Q$ in the domain, there is a single turning point. Denote the zero of $Q$ by $\mu$. Then, in this case, the solutions are exponential in the region $1\leq s\leq\mu$, and oscillatory in the region $s_{0}\leq s\leq r_{2}/r_{1}$. Using the technique of Langer (1959) we can approximate the solution in this case as
@@ A.2 Two turning point case
When $V_{A}\ll 1$, there are two roots of $Q$ located at $\left(\frac{m}{2\omega}\left(1+\sqrt{1\pm 4\omega d_{i}V_{A}\sqrt{m^{2}-1/4}}\right)\right)^{2/3}$. In this case, the domain is divided into two regions where the solution is wavelike, $Q>0$ separated by an ‘evanescent region’ region where the solution is damped, $Q<0$. Let $\mu_{1},\mu_{2}$ denote the locations of the two zeroes of $Q$, and assume $Q<0$ for $\mu_{1}\leq s\leq\mu_{2}$. Then, it can be shown using a Liouville-Green transformation that the solutions are asymptotic to parabolic cylinder functions  (Nayfeh, 2000). In particular, define $\zeta(s)$ such that when $\mu_{1}\leq s\leq\mu_{2}$
@@
$a$ can be interpreted as the size of the ‘classically forbidden region’ in a quantum mechanical tunneling problem. Similar expressions involving suitably chosen branches of the square roots ensure $\zeta(s)$ is a regular function, and that the regions where $Q>0$ are mapped to $\zeta^{2}>1$, and the region where $Q<0$ is mapped to the region $\zeta^{2}<1$. Namely, for $1\leq s\leq\mu_{1}$, set
@@
where $E,E^{*}$ are the complex solutions given by Abramowitz and Stegun (1974). As noted by Narayan et al. (1987), the functions $E$ and $E^{*}$ represent radially outward and inward propagating waves, respectively. The boundary conditions on $u$ are transformed into the conditions